# Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning

### v1

@@ Abstract
Deep neural network models have achieved remarkable progress in 3D scene understanding while trained in the closed-set setting and with full labels. However, the major bottleneck for the current 3D recognition approach is that these models do not have the capacity to recognize any unseen novel classes beyond the training categories in diverse kinds of real-world applications. In the meantime, current state-of-the-art 3D scene understanding approaches primarily require a large number of high-quality labels to train neural networks, which merely perform well in a fully supervised manner. Therefore, we are in urgent need of a framework that can simultaneously be applicable to both 3D point cloud segmentation and detection, particularly in the circumstances where the labels are rather scarce. This work presents a generalized and simple framework for dealing with 3D scene understanding when the labeled scenes are quite limited. To extract knowledge for novel categories from the pre-trained vision-language models, we propose a hierarchical feature-aligned pre-training and knowledge distillation strategy to extract and distill meaningful information from large-scale vision-language models, which helps benefit the open-vocabulary scene understanding tasks. To leverage the boundary information, we propose a novel energy-based loss with boundary awareness benefiting from the region-level boundary predictions. To encourage latent instance discrimination and to guarantee efficiency, we propose the unsupervised region-level semantic contrastive learning scheme for point clouds, using confident predictions of the neural network to discriminate the intermediate feature embeddings at multiple stages. In the limited reconstruction case, our proposed approach, termed WS3D++, ranks 1st on the large-scale ScanNet benchmark on both the task of semantic segmentation and instance segmentation. Also, our proposed WS3D++ achieves state-of-the-art data-efficient learning performance on the other large-scale real-scene indoor and outdoor datasets S3DIS and SemanticKITTI. Extensive experiments with both indoor and outdoor scenes demonstrated the effectiveness of our approach in both data-efficient learning and open-world few-shot learning. All codes, models, and data are to made publicly available at: https://github.com/KangchengLiu. The code is at: WS3D++ Code link.
@@ 1 Introduction
The 3D scene parsing problem, which typically encompasses several important downstream tasks: point cloud semantic segmentation, instance segmentation, and object detection, becomes increasingly important with the wide deployment of 3D sensors, such as LiDAR and RGB-D cameras [1, 2, 3, 4, 5, 6, 7, 8]. Point clouds are raw sensor data obtained from 3D sensors and the most simple and common 3D data representation for understanding 3D scenes of robot navigation, robot grasping, and manipulation tasks. Despite significant success in deep neural networks applied to 3D visual perception, two major challenges hinder the construction of more scalable visual perception systems in 3D worlds. One is the closed-set assumption, which means the model only performs well while recognizing the categories that appear in the training set and struggles in recognizing the novel unseen object categories or concepts. Another is the heavy reliance on large amounts of high-quality labeled data. Large-scale 3D scenes are very laborious to label, which also makes it very hard for deep network models to perform well with very limited annotations.
@@
Reliance on large-scale labeled data. A critical prerequisite for fully exploiting the capacity of the fully supervised deep learning approaches is the accessibility to large-scale well-annotated high-quality training data. Most point cloud understanding methods rely on heavy annotations [17, 18, 19]. However, the annotation of large-scale 3D point cloud scenes is rather time-consuming and labor-intensive. For instance, it requires around thirty minutes to label a single scene for ScanNet [20] or S3DIS [21] with thousands of scenes. Though existing point cloud understanding methods [17, 18, 19] have achieved decent results on these datasets, it is difficult to directly extend them to novel scenes when the high-quality labeled data is scarce. In the meanwhile, it is often the case that a limited number of scenes can be reconstructed in real applications [22]. Therefore, developing methods that can be trained with very limited labeled scenes, termed data-efficient 3D scene understanding with limited scene-level annotation, becomes in high demand. Data-efficient semantic and instance segmentation [23, 24, 25] is a vehement research topic for image-level scene understanding. Some simple but successful methods have been proposed, such as contrastive learning [26, 27] which learns a meaningful and discriminative representation, and conditional random field (CRF) [28, 29] for pseudo label propagation. However, there still exist four main challenging unsolved issues while scaling up these approaches to 3D scene understanding. First, the widely adopted energy function-based conditional random field segmentation [29] relies on handcrafted feature similarities and does not consider explicit boundary information. It attaches equal importance to pixels on semantic boundaries and within same semantic objects, which can cause vague and inaccurate predictions in pixel-level segmentation at object boundaries. And how to leverage boundary information has been explored in 2D but rarely explored in 3D data efficient learning [30]. Second, the computation costs are both very high when applying point-level contrastive learning or point-level energy-based segmentation in a dense point cloud scene for every point pair [31]. Furthermore, large-scale point cloud scenes even contain billions of points, making point-level contrastive learning intractable in computational costs. Third, the existing unsupervised contrastive learning-based pre-training for point clouds [27, 22, 32, 33] only considers geometrically registered point/voxel pairs as the positive samples, while it does not explicitly consider explicit regional information, let alone the hierarchical alignments.
Figure 1: Overall illustrative diagram of our proposed WS3D++. We integrate language-3D feature associated pre-training and data-efficient fine-tuning to achieve effective data-efficient and open-vocabulary 3D scene understanding for 3D scenes.
@@
Driven by the above motivations in terms of both generalization capacity and data efficiency, we propose an effective two-stage framework, involving unsupervised hierarchical vision-language pre-training and label-efficient fine-tuning to enable more label/data-efficient 3D scene understanding. As shown in Fig. 1, in the pre-training stage, we leverage the rendering techniques to construct well-aligned 2D views for large-scale 3D scenes to establish more accurate coarse-to-fine vision-language associations. Then, we leverage the off-the-shelf object detector and the pre-trained large-scale vision-language model CLIP [9] to construct the hierarchical feature representations from both the global scene level to the local object level. We also propose an effective knowledge distillation strategy that distills the informative visual-language-aligned representation of the image encoder in CLIP [9] to the 3D backbone network.
@@
First, we propose a generalized pre-training approach for data-efficient learning, which establishes accurate alignments between language and 3D point cloud in terms of both object-level and scene-level semantics in a hierarchical manner.  Second, we propose to leverage rendering techniques which make explicit associations between image and point cloud to facilitate 2D-to-3D matching and subsequent language-to-3D matching.  Third, we visualized the language-queried activation maps directly on the 3D scenes, which demonstrates that the proposed approach learns better visual-linguistic alignment between the language descriptions and the visual object-level information. Finally, we evaluate our proposed approach comprehensively in diverse data-efficient and open-world learning settings for both the 3D semantic segmentation and 3D instance segmentation tasks.
@@
    During the pre-training stage, we first design an effective approach which distills rich knowledge from the large-scale vision-language model into the 3D point cloud modality. Specifically, we propose leveraging rendering to obtain explicit scene-level and object-level 2D-3D feature associations, establishing a more accurate vision-language association hierarchically than the original CLIP encoder.
@@
    During the pre-training stage, we further propose a word-to-3D matching approach to establish the well-aligned language-3D associated feature representations at both the scene level and the object level, which largely facilitates the subsequent effective contrastive learning with the mostly matched visual-linguistic contrastive pairs. The proposed designs have both enhanced the data-efficient learning and the knowledge-transfer capacity of the model, as demonstrated by our extensive experiments on both the 3D object detection and the 3D semantic/instance segmentation tasks.
@@
To the best of our knowledge, this is the pioneer work which comprehensively evaluates across diverse 3D label-efficient scene understanding downstream tasks with our proposed 3D open-vocabulary recognition approach termed WS3D++. Our endeavor is orthogonal to the 3D backbone network designs and thus can be seamlessly integrated with the prevailing 3D point cloud detection or segmentation models. Our comprehensive results provide solid baselines for future research in data-efficient 3D scene understanding.
@@ 2 Related Work
Learning-based Point Cloud Understanding. Deep-network-based approaches are widely adopted for point cloud understanding. Fully supervised approaches can be roughly categorized into voxel-based [36, 37, 38, 39, 40], projection-based [41, 42, 43, 44, 45, 46], and point-based approaches [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. The voxel-based approaches [59, 60, 61] which are built upon SparseConv [59] and voxelize the point cloud for efficient processing have achieved remarkable performance in 3D scene parsing. Therefore, we use SparseConv [59] as our backbone architecture for downstream semantic understanding tasks because of its high performance in inferring 3D semantics. Also, the LiDAR-based approaches are of significance to many industrial applications such as inspections [62, 63, 64] and robot localization [65, 66, 67, 68, 69], and large-scale robotic scene parsing [57, 70, 71], as well as robotic control applications [72, 73, 74, 75, 76, 77], etc.
@@
Recently, with the development of large-scale vision-language models such as CLIP [9] and Flamingo [10], we can largely benefit the recognition capacity of 3D scene understanding models by distilling knowledge from large-scale vision-language models. For example, some pioneer works such as PointCLIP [82, 83] and ULIP [84] successfully transfer the knowledge from vision-language models to boost the downstream 3D shape classification tasks. The 3D CAD shapes are transferred to multi-view depth maps, thus they can be fed into the CLIP visual encoder and used the image representations paired with the corresponding point cloud as a bridge to obtain the correlations between the 3D and textual features. Moving beyond object-level recognition tasks, pioneer works also explore how to establish alignment among images, language, and 3D point cloud scenes for the task of open-vocabulary 3D scene understanding [85, 86, 87], which target localizing, detecting, and segmenting novel object categories that do not exist in the annotation. Compared with them, our proposed simple but effective framework can be both applicable to data-efficient learning and open-vocabulary scene understanding.
Figure 2: The pre-training paradigm of our proposed WS3D++. We propose hierarchical feature alignments to establish hierarchical vision-language aligned feature representations during pre-training. This kind of paradigm helps learn more powerful visual-linguistic matched representations ranging from both the global view-level to the local object category-level.
@@
Label/Data-Efficient Learning for 3D. Recent studies have produced many elaborately designed backbone networks for 3D semantic/instance segmentation [88, 89, 45, 90, 91], as well as for 3D object detection [92, 93, 94, 95]. However, they rely on full supervision. Directly applying current SOTAs (State-of-the-art) methods for training will result in a great decrease in performance [96] for WSL, if the percentage of labeled data drops to a certain value, e.g., less than 30%. Recently, many works have started to focus on point cloud semantic segmentation with partially labeled data. Wang et al. [97] choose to transform point clouds to images, but pixel-level semantic segmentation labels are required for network training. Sub-cloud annotations [98] require extra labor to separate the sub-clouds and to label points within the sub-clouds. Liu [99] proposed a robust data-efficient 3D scene parsing framework. It leverages the complementary merits of the superior generalization capacity of the traditional 3D descriptors and the strong feature description capacity learned 3D descriptors to learn very robust local features. Then using the descriptor guided learned region merging [3], superior performance can be achieved on downstream tasks. Liu et al. [100, 101] proposed self-training techniques to tackle scene understanding in weak supervision, which design a two-stage training scheme to produce iteratively optimized pseudo labels from weak labels during training. Despite satisfactory results, these approaches still have not learned generalized representation applicable to diverse tasks. Xu et al. [102] adopt a weakly-supervised training strategy, which combines training with coarse-grained information and partial points using 10% labels. However, their tested cases are limited to object part segmentation, and it is difficult to uniformly choose points to label. The convex decomposition [103] is conducted in an approximate manner to perform 3D scene parsing on the object parts. More approaches [104] have been proposed recently, which utilize class prototypes and masked point cloud modeling [105, 106, 81] to learn informative representations for downstream 3D scene understanding. To sum up, although approaches have been proposed to alleviate the data efficiency problem, the models for weakly supervised learning lack the capacity to recognize novel categories beyond the labeled training set. Our framework tackles both open-set and data-efficient learning problems and is widely applicable to diverse 3D scene understanding downstream tasks.

### v2

# Generalized Robot 3D Vision-Language Model with Fast Rendering and Pre-Training Vision-Language Alignment

@@ Abstract
Deep neural network models have achieved remarkable progress in 3D scene understanding while trained in the closed-set setting and with full labels. However, the major bottleneck for the current 3D recognition approach is that these models do not have the capacity to recognize any unseen novel classes beyond the training categories in diverse real-world applications. In the meantime, current state-of-the-art 3D scene understanding approaches primarily require a large number of high-quality labels to train neural networks, which merely perform well in a fully supervised manner. Therefore, we are in urgent need of a framework that can simultaneously be applicable to both 3D point cloud segmentation and detection, particularly in the circumstances where the labels are rather scarce. This work presents a generalized and straightforward framework for dealing with 3D scene understanding when the labeled scenes are quite limited. To extract knowledge for novel categories from the pre-trained vision-language models, we propose a hierarchical feature-aligned pre-training and knowledge distillation strategy to extract and distill meaningful information from large-scale vision-language models, which helps benefit the open-vocabulary scene understanding tasks. To leverage the boundary information, we propose a novel energy-based loss with boundary awareness benefiting from the region-level boundary predictions. To encourage latent instance discrimination and to guarantee efficiency, we propose the unsupervised region-level semantic contrastive learning scheme for point clouds, using confident predictions of the neural network to discriminate the intermediate feature embeddings at multiple stages. In the limited reconstruction case, our proposed approach, termed WS3D++, ranks 1st on the large-scale ScanNet benchmark on both the task of semantic segmentation and instance segmentation. Also, our proposed WS3D++ achieves state-of-the-art data-efficient learning performance on the other large-scale real-scene indoor and outdoor datasets S3DIS and SemanticKITTI. Extensive experiments with both indoor and outdoor scenes demonstrated the effectiveness of our approach in both data-efficient learning and open-world few-shot learning. The code is at: WS3D++ Code link.
@@ 1 Introduction
Figure 1: The final overall illustrative diagram of our proposed WS3D++. We integrate language-3D feature associated pre-training and data-efficient fine-tuning as a general scene parsing vision-language model to achieve effective data-efficient as well as open-vocabulary 3D scene understanding for 3D scenes.
The 3D scene parsing problem, which typically encompasses several important downstream tasks: point cloud semantic segmentation, instance segmentation, and object detection, becomes increasingly important with the wide deployment of 3D sensors, such as LiDAR and RGB-D cameras [1, 2, 3, 4, 5, 6]. Point clouds are raw sensor data obtained from 3D sensors and the most simple and common 3D data representation for understanding 3D scenes of robot navigation, robot grasping, and manipulation tasks. Despite significant success in deep neural networks applied to 3D visual perception, two major challenges hinder the construction of more scalable visual perception systems in 3D worlds. One is the closed-set assumption, which means the model only performs well while recognizing the categories that appear in the training set and struggles in recognizing the novel unseen object categories or concepts. Another is the heavy reliance on large amounts of high-quality labeled data. Large-scale 3D scenes are very laborious to label, which also makes it very hard for deep network models to perform well with very limited annotations.
Figure 2: The pre-training paradigm of our proposed WS3D++. We propose the hierarchical global to local feature alignments to establish the hierarchical vision-language aligned feature representations during the pre-training. This proposed paradigm helps to learn more powerful visual-linguistic aligned feature representation during the pre-training stage. We have further shown the final visualizations and comparisons with CLIP text presentations ranging from both the global view level to the local object category level. The results have further demonstrate the vision-language aligned feature representation for 3D scene parsing.
@@
Figure 3: The feature matching visualization of our proposed WS3D++. We propose hierarchical global to local feature alignments to establish hierarchical vision-language aligned feature representations during pre-training from both the global view level to the local object level. This kind of paradigm helps to learn more powerful visual-linguistic matched representations ranging from both the global view-level to the local object category-level. In the above figure, we have shown the matching at the global view on the left and the matching at local object level on the right. It can be demonstrated that our proposed approach can establish matched feature representation at both the global room feature level and the local object feature level.
Figure 4: The fine-tuning paradigm of our proposed WS3D++. WS3D++ [4] consists of three proposed modules: 1. The unsupervised region-level energy-based optimization guided by boundary labels; 2. The unsupervised multi-stage region-level contrastive learning with high confidence; 3. The supervised region-level semantic contrastive learning with labeled data. The backbone network adopts encoder-decoder structures. The weights of the backbone network are shared in the supervised and unsupervised branches. Integrated with the proposed pre-training paradigm illustrated in Figure 2, by our proposed hierarchical feature aligned pre-training and regional fine-tuning, more effective label-efficient learning as well as open-vocabulary learning is realized.
@@
Reliance on large-scale labeled data. A critical prerequisite for fully exploiting the capacity of the fully supervised deep learning approaches is the accessibility to large-scale well-annotated high-quality training data. Most point cloud understanding methods rely on heavy annotations [15, 16, 17]. However, the annotation of large-scale 3D point cloud scenes is rather time-consuming and labor-intensive. For instance, it requires around thirty minutes to label a single scene for ScanNet [18] or S3DIS [19] with thousands of scenes. Though existing point cloud understanding methods [15, 16, 17] have achieved decent results on these datasets, it is difficult to directly extend them to novel scenes when the high-quality labeled data is scarce. In the meanwhile, it is often the case that a limited number of scenes can be reconstructed in real applications [20]. Therefore, developing methods that can be trained with very limited labeled scenes, termed data-efficient 3D scene understanding with limited scene-level annotation, becomes in high demand. Data-efficient semantic and instance segmentation [21, 22, 23] is a vehement research topic for image-level scene understanding. Some simple but successful methods have been proposed, such as contrastive learning [24, 25] which learns a meaningful and discriminative representation, and conditional random field (CRF) [26, 27] for pseudo label propagation. However, there still exist four main challenging unsolved issues while scaling up these approaches to 3D scene understanding. First, the widely adopted energy function-based conditional random field segmentation [27] relies on handcrafted feature similarities and does not consider explicit boundary information. It attaches equal importance to pixels on semantic boundaries and within same semantic objects, which can cause vague and inaccurate predictions in pixel-level segmentation at object boundaries. And how to leverage boundary information has been explored in 2D but rarely explored in 3D data efficient learning [28]. Second, the computation costs are both very high when applying point-level contrastive learning or point-level energy-based segmentation in a dense point cloud scene for every point pair [29, 30]. Furthermore, large-scale point cloud scenes even contain billions of points, making point-level contrastive learning intractable in computational costs. Third, the existing unsupervised contrastive learning-based pre-training for point clouds [25, 20, 31, 32] only considers geometrically registered point/voxel pairs as the positive samples, while it does not explicitly consider explicit regional information, let alone the hierarchical alignments.
@@
Driven by the above motivations in terms of both generalization capacity and data efficiency, we propose an effective two-stage framework, involving unsupervised hierarchical vision-language pre-training and label-efficient fine-tuning to boost the label/data-efficiency in 3D scene understanding. As shown in Fig. 1, in the pre-training stage, we leverage the rendering techniques to construct well-aligned 2D views for large-scale 3D scenes to establish more accurate coarse-to-fine vision-language associations. Then, we leverage the off-the-shelf object detector and the pre-trained large-scale vision-language model CLIP [7] to construct the hierarchical feature representations from both the global scene level to the local object level. We also propose an effective knowledge distillation strategy that distills the informative visual-language-aligned representation of the image encoder in CLIP [7] to the 3D backbone network. As is demonstrated by our extensive experiments, the open-vocabulary scene parsing performance can be also enhanced.
@@
First, we propose a generalized pre-training approach for data-efficient learning, which establishes accurate alignments between language and 3D point cloud in terms of both object-level and scene-level semantics in a hierarchical manner.  Second, we propose to leverage the rendering technique that makes explicit associations between image and point cloud to facilitate 2D-to-3D matching and subsequent language-to-3D matching.  Third, we visualized the language-queried activation maps directly on the 3D scenes, which demonstrates that the proposed approach learns better visual-linguistic alignment between the language descriptions and the visual object-level information. Finally, we evaluate our proposed approach comprehensively in diverse data-efficient and open-world learning settings for both the 3D semantic segmentation and 3D instance segmentation tasks.
@@
    During the pre-training stage, we first propose an effective design which distills rich knowledge from the large-scale vision-language model into the 3D point cloud modality. Specifically, we propose leveraging rendering to obtain explicit scene-level and object-level 2D-3D feature associations, establishing a more accurate vision-language association hierarchically than the original CLIP encoder. We have demonstrated by extensive experiments that our proposed approach can realize superior compatibility with prevailing weakly supervised approaches.
@@
    During the pre-training stage, we first propose a global scene-to-sentence matching and then propose a local object-to-word matching approach, respectively, to establish the well-aligned vision-language feature representations at both the scene level and the object level, which largely facilitates the subsequent effective contrastive learning with the mostly matched visual-language contrastive pairs. The proposed designs have both enhanced the data-efficient learning and the knowledge-transfer capacity of the model, as demonstrated by our extensive experiments on both the 3D object detection and the 3D semantic/instance segmentation tasks.
@@
    Figure 5: Qualitative semantic segmentation results of the proposed WS3D++ on SemanticKITTI validation set the with the 5% labeling percentage, compared with the fully supervised state-of-the-art Cylinder3D [34], and BAAF-Net [35] with the diverse semantics indicated by different colors. The red circles highlight the final performance difference between diverse comparative approaches.
@@
To the best of our knowledge, this is the pioneer work which comprehensively evaluates across diverse 3D label-efficient scene understanding downstream tasks with our proposed 3D open-vocabulary recognition approach termed WS3D++. Our endeavor is orthogonal to the 3D backbone network designs and thus can be seamlessly integrated with the prevailing 3D point cloud detection or segmentation models. Our comprehensive results provide solid baselines for future researches in the data-efficient 3D scene understanding.
Figure 6: The captured 2D and 3D region proposals. It is demonstrated qualitatively clearly that more precise object proposals are captured by proposed united 2D/3D proposal generation approach. It can be demonstrated that clear superior regional proposal generation performance can be well guaranteed.
Figure 7: The final ScanNet [18] object detection results and performance. The comparative differences in detection predictions are highlighted in the rectangles.
@@ 2 Related Work
Learning-based Point Cloud Understanding. Deep-network-based approaches are widely adopted for point cloud understanding. Fully supervised approaches can be roughly categorized into voxel-based [37, 38, 39, 40, 41], projection-based [42, 43, 44, 45, 46, 47], and point-based approaches [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. The voxel-based approaches [59, 60, 61] which are built upon SparseConv [59] and voxelize the point cloud for efficient processing have achieved remarkable performance in 3D scene parsing. Therefore, we use SparseConv [59] as our backbone architecture for downstream semantic understanding tasks because of its high performance in inferring 3D semantics.
@@
Recently, with the development of large-scale vision-language models such as CLIP [7] and Flamingo [8], we can largely benefit the recognition capacity of 3D scene understanding models by distilling informative knowledge from large-scale vision-language models. For example, some pioneer works such as PointCLIP [66, 67] and ULIP [68] successfully transfer the knowledge from vision-language models to boost the downstream 3D shape classification tasks. The 3D CAD shapes are transferred to multi-view depth maps, thus they can be fed into the CLIP visual encoder and used the image representations paired with the corresponding point cloud as a bridge to obtain the correlations between the 3D and textual features. Moving beyond object-level recognition tasks, pioneer works also explore how to establish alignment among images, language, and 3D point cloud scenes for the task of open-vocabulary 3D scene understanding [69, 70, 71], which target localizing, detecting, and segmenting novel object categories that do not exist in the annotation. Compared with them, our proposed simple but effective framework can be both applicable to data-efficient learning and open-vocabulary scene understanding.
@@
Label/Data-Efficient Learning for 3D. Recent studies have produced many elaborately designed backbone networks for 3D semantic/instance segmentation [72, 35, 46, 73, 74], as well as for 3D object detection [75, 76, 77]. However, they rely on full supervision. Directly applying current SOTAs (State-of-the-art) methods for training will result in a great decrease in performance [78] for WSL, if the percentage of labeled data drops to a certain value, e.g., less than 30%. Recently, many works have started to focus on point cloud semantic segmentation with partially labeled data. Wang et al. [79] choose to transform point clouds to images, but pixel-level semantic segmentation labels are required for network training. Sub-cloud annotations [80] require extra labor to separate the sub-clouds and to label points within the sub-clouds. Liu [81] proposed a robust data-efficient 3D scene parsing framework. It leverages the complementary merits of the superior generalization capacity of the traditional 3D descriptors and the strong feature description capacity learned 3D descriptors to learn very robust local features. Then using the descriptor guided learned region merging [3], superior performance can be achieved on downstream tasks. Liu et al. [82, 83] proposed self-training techniques to tackle scene understanding in weak supervision, which design a two-stage training scheme to produce iteratively optimized pseudo labels from weak labels during training. Despite satisfactory results, these approaches still have not learned generalized representation applicable to diverse tasks. Xu et al. [84] adopt a weakly-supervised training strategy, which combines training with coarse-grained information and partial points using 10% labels. However, their tested cases are limited to object part segmentation, and it is difficult to uniformly choose points to label. The convex decomposition [85] is conducted in an approximate manner to perform 3D scene parsing on the object parts. More approaches [86] have been proposed recently, which utilize class prototypes and masked point cloud modeling [87, 88, 65] to learn informative representations for downstream 3D scene understanding. Conceptfusion [89] has also been proposed to alleviate the labelling burden via rendering with the proposed CLIP-driven queryable 3D point cloud maps. To sum up, although approaches have been proposed to alleviate the data efficiency problem, the models for weakly supervised learning lack the capacity to recognize novel categories beyond the labeled training set. Our framework tackles open-set and data-efficient learning problems and is widely applicable to diverse 3D scene understanding downstream tasks.
TABLE I: Comparisons of the open-vocabulary learning performance on ScanNet. It can be demonstrated that our proposed approach provides very superior open-world recognition performance compared with the diverse SOTAs. The results are given as hIoU / mIoUB / mIoUN, respectively. We have compared intensively with diverse labeled ratios. We apply panoptic quality mean intersection over union (mIoU), which can be divided into segmentation quality and recognition quality as metrics for the instance segmentation task. These evaluation metrics are computed on base (B) and novel (N) categories, with the superscripts of B and N (e.g. mIoUB / mIoUN), respectively. Furthermore, we have also utilized the harmonic metric such as harmonic IoU (hIoU) as the major indicators for open-world tasks following popular zero-shot learning work to consider the partition between base and novel classes.
Datasets Models 80% 100%
B15/N4 B12/N7 B10/N9 B15/N4 B12/N7 B10/N9
ScanNet [18] Baseline 39.7 / 48.8 / 38.9 37.9 / 44.9 / 42.6 31.9 / 38.8 / 33.9 43.6 / 50.7 / 39.6 38.8 / 48.9 / 45.8 33.9 / 42.9 / 38.8
Merely the Pre-training Stage 55.9 / 52.9 / 48.8 46.9 / 48.8 / 46.9 43.9 / 46.7 / 43.8 69.6 / 67.8 / 59.8 55.8 / 50.9 / 48.8 53.6 / 47.9 / 46.8
3DGenZ [90] 32.6 / 76.8 / 26.6 25.9 / 46.5 / 26.3 28.3 / 63.6 / 18.6 42.9 / 76.9 / 39.6 32.9 / 56.8 / 33.9 27.1 / 69.8 / 32.8
3DGenZ [90] (Pr) 42.8 / 78.8 / 42.9 39.6 / 68.3 / 51.8 27.6 / 79.5 / 18.9 46.8 / 88.6 / 48.6 43.8 / 78.8 / 49.9 37.9 / 83.7 / 38.9
3DTZSL [91] 17.6 / 45.8 / 6.1 23.8 / 36.6 / 12.8 13.8 / 67.5 / 15.7 42.9 / 49.6 / 16.7 27.8 / 38.9 / 15.9 16.9 / 69.8 / 18.8
3DTZSL [91] (Pr) 49.6 / 58.7 / 48.9 38.8 / 50.6 / 38.6 33.9 / 69.8 / 33.9 59.6 / 69.8 / 59.6 48.9 / 59.6 / 48.7 38.6 / 73.9 / 39.6
LSeg3D [92] 0.0 / 64.4 / 0.0 26.9 / 55.7 / 22.8 11.8 / 77.6 / 23.9 33.6 / 76.8 / 33.8 28.8 / 72.6 / 29.2 26.6 / 68.5 / 26.9
LSeg3D [92] (Pr) 38.9 / 78.8 / 29.9 32.8 / 73.6 / 26.8 28.6 / 66.8 / 22.8 49.6 / 82.7 / 46.9 46.9 / 76.8 / 43.2 38.9 / 70.3 / 39.6
PLA without caption [69] 39.7 / 73.3 / 28.0 31.5 / 72.3 / 25.6 28.7 / 68.6 / 23.5 45.7 / 72.9 / 36.9 31.9 / 77.8 / 26.8 29.8 / 79.8 / 28.3
PLA without caption [69] (Pr) 53.9 / 82.6 / 45.8 48.6 / 80.6 / 40.8 42.6 / 75.5 / 37.8 56.7 / 84.3 / 52.9 53.9 / 83.8 / 45.3 52.1 / 80.8 / 45.1
PLA [69] 69.2 / 70.6 / 66.9 60.9 / 68.9 / 59.9 58.2 / 78.6 / 73.8 69.3 / 73.3 / 72.2 62.7 / 73.6 / 61.3 65.6 / 81.8 / 76.8
PLA (Pr) [69] 79.9 / 82.9 / 81.9 78.5 / 80.6 / 68.9 65.9 / 77.8 / 57.8 82.5 / 88.2 / 84.9 80.8 / 85.7 / 73.9 76.8 / 85.3 / 63.9
WS3D++ (Ours) 72.3 / 69.1 / 73.3 68.6 / 65.6 / 68.6 73.2 / 71.9 / 71.6 77.9 / 76.6 / 78.6 73.8 / 72.8 / 75.8 71.9 / 72.8 / 73.9
WS3D++ (Ours) (Pr) 75.8 / 73.8 / 77.5 78.6 / 75.3 / 78.9 71.8 / 85.6 / 79.5 86.7 / 83.7 / 85.8 82.7 / 80.9 / 83.6 78.9 / 88.9 / 83.8
Datasets Models 10% 50%
B15/N4 B12/N7 B10/N9 B15/N4 B12/N7 B10/N9
ScanNet [18] Baseline 27.9 / 18.2 / 21.9 23.0 / 33.6 / 31.2 19.6 / 28.6 / 27.6 35.1 / 20.6 / 36.6 31.7 / 41.5 / 38.2 28.5 / 32.3 / 28.2
Merely the Pre-training Stage 35.8 / 58.3 / 36.8 33.8 / 37.9 / 31.8 30.8 / 70.3 / 25.6 37.2 / 63.2 / 39.7 48.9 / 53.8 / 49.9 49.6 / 86.8 / 49.8
3DGenZ [90] 17.3 / 52.3 / 9.7 15.3 / 28.8 / 12.5 11.2 / 62.5 / 5.2 18.6 / 52.8 / 11.6 17.8 / 33.6 / 12.9 11.8 / 72.3 / 6.3
3DGenZ [90] (Pr) 28.5 / 72.9 / 28.2 23.6 / 38.9 / 23.6 15.6 / 68.9 / 12.9 29.7 / 75.6 / 29.7 29.5 / 45.3 / 45.7 23.3 / 43.9 / 11.9
3DTZSL [91] 8.9 / 33.9 / 6.1 3.3 / 35.8 / 1.8 7.3 / 47.3 / 2.8 16.2 / 32.5 / 23.2 21.9 / 31.3 / 6.8 9.9 / 63.5 / 8.8
3DTZSL [91] (Pr) 26.8 / 43.8 / 28.7 18.6 / 52.6 / 23.2 16.8 / 49.5 / 19.2 31.3 / 49.3 / 28.7 28.6 / 43.9 / 27.3 15.3 / 66.9 / 21.7
LSeg3D [92] 0.0 / 55.6 / 0.0 0.6 / 43.6 / 0.3 1.2 / 56.1 / 0.8 19.9 / 65.3 / 13.6 12.9 / 56.8 / 8.7 9.9 / 58.5 / 8.7
LSeg3D [92] (Pr) 28.8 / 65.3 / 23.8 26.7 / 62.8 / 16.7 23.7 / 62.8 / 22.9 32.8 / 76.7 / 28.7 28.9 / 73.9 / 26.9 26.8 / 73.2 / 25.6
PLA without caption [69] 35.3 / 53.2 / 18.2 26.7 / 52.1 / 9.9 19.6 / 56.7 / 9.7 36.5 / 55.9 / 22.8 28.2 / 58.7 / 18.8 25.8 / 62.1 / 15.2
PLA without caption [69] (Pr) 39.3 / 73.3 / 35.7 28.6 / 67.8 / 29.6 24.5 / 62.6 / 26.8 42.9 / 75.6 / 41.8 39.3 / 73.6 / 36.9 35.6 / 72.3 / 33.8
PLA [69] 51.3 / 58.8 / 48.6 49.3 / 55.9 / 44.9 44.7 / 52.2 / 40.8 65.7 / 72.8 / 65.7 58.8 / 66.5 / 49.8 53.6 / 63.5 / 46.8
PLA (Pr) [69] 61.8 / 67.6 / 73.9 58.7 / 59.3 / 56.8 68.8 / 82.9 / 47.6 72.5 / 82.8 / 79.9 71.8 / 81.9 / 61.8 70.3 / 83.8 / 49.8
WS3D++ (Ours) 65.9 / 75.2 / 68.3 61.3 / 65.6 / 62.7 58.6 / 60.9 / 56.8 71.8 / 76.8 / 81.6 65.8 / 69.5 / 68.9 62.8 / 67.8 / 68.2
WS3D++ (Ours) (Pr) 69.9 / 68.8 / 67.8 67.9 / 67.7 / 66.9 63.6 / 66.2 / 65.6 70.9 / 70.8 / 71.9 72.8 / 71.8 / 69.8 69.8 / 75.6 / 72.1